OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (y – 1) = 3(x + 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

C by the power invested in me

OpenStudy (whpalmer4):

Point slope form for a line with slope \(m\) passing through point \((x_1,y_1)\) is \[y-y_1 = m(x-x_1)\]

OpenStudy (whpalmer4):

Part of the problem with just taking answers posted without any explanation is that you have little way of knowing if they were posted by someone who knew how to do the work correctly.

OpenStudy (whpalmer4):

I suggest that you be skeptical about the powers invested in @mattmiller :-)

OpenStudy (anonymous):

so it is c right thats the answer i got i did it twice i got c

OpenStudy (whpalmer4):

Plug in the values x = 2 and y = -1 in your answer. Does the result make sense?

OpenStudy (whpalmer4):

\[(x+2) = 3(y-1)\]\[x=2\]\[y=-1\]\[(2+2) = 3(-1-1)\]\[4 = 3(-2)\]\[4=-6\]Not in my experience.

OpenStudy (mathmale):

@OOOOOOOO: I really think it'd be faster, and help you a lot more, if you'd learn how to find the equation of lines in point-slope form (and in slope-intercept form) yourself. This guessing game is so sterile.

OpenStudy (mathmale):

whpalmer4's equation, \[y-y_1 = m(x-x_1)\]

OpenStudy (mathmale):

along with the slope m=3, and point (2,-1), are all you need to work out this problem, by substitution. Please go for it; share y our work; ask questions if you need to.